solve the IVP using the annihilator approach

y(3) + 9y' = excos(3x)

y(0) = 2

y'(0) = 1

y''(0) = 1
Answer #1

Given equation is. y + gy! =e* cos(3x) let D= d dt then y = D3 y and - DY ou D3y+9DY - SO, equation 18. ex Cosc3x) → (D3 + 9so DP = -9. Yp e 22 Cos(32) 11 121 - 28 3D +7 een 4 Cos (3x) 9 DP-49 e 2x (3D + 7) cos(3x) DE-9 520 2x 520 e² [-9 sin(3x) + 7using initial conditions. yo) - C +2 + 7 520 - 2 y(o) 3C3 + = 27 - .14 520 = 1 S20 y (0) -962 14 11 BI S20 520 we hare. 103